Grade 12th passAlgebra

Lakshmi is a cashier in a bank. She has currency notes of denominations ₹100, ₹50 and ₹10, respectively. The ratio of the number of these notes is 2:3:5. The total cash with Lakshmi is ₹4,00,000. How many notes of each denomination does she have?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the numbers of notes of ₹100, ₹50 and ₹10 be 2x, 3x and 5x respectively. Valueof₹100=2x×100=200x
Valueof₹50=3x×50=150x
Value of ₹10 = 5x × 10 = 50x
According to the question,
200x + 150x + 50x = 4,00,000
⇒ 400x = 4,00,000
⇒x = 400000/400
⇒x=1000
Numbers of ₹100 notes = 2x = 2000
Numbers of ₹50 notes = 3x = 3000
Numbers of ₹10notes = 5x = 5000

Thanks